# Starfish lemma via birational quasi-isomorphisms

### v2

@@ 4.2 Transferring cluster and algebraic properties
In this subsection, we show how to derive the conditions of the Starfish lemma for $(\mathcal{C},R)$ via $\mathcal{Q}$ given they are satisfied for $(\tilde{\mathcal{C}},\tilde{R})$ and the marked variables. The main result is Proposition 4.9.
@@
  2. 2)
    Both $\tilde{x}$ and its mutation $\tilde{x}^{\prime}$ are regular and coprime.
@@
For the next proposition, let $\mathbf{x}_{i}^{\prime}$ denote the mutation of $\mathbf{x}_{0}$ in direction $i\in[1,N]$.
@@
Assume that the initial exchange matrices in $\mathcal{C}$ and $\tilde{\mathcal{C}}$ are of full rank, as well as $\bar{\mathcal{A}}(\tilde{\mathcal{C}})\supseteq\tilde{R}$, and for each $i\in\mathcal{I}(\kappa)$, $\mathcal{L}(\mathbf{x}_{i}^{\prime})\supseteq R$. Then $\bar{\mathcal{A}}(\mathcal{C})\supseteq R$.
@@
4.3 Starfish lemma
In this subsection, we prove our first main result on how to derive the conditions of the Starfish lemma for $(\mathcal{C},R)$ via $\mathcal{Q}$ given they are satisfied for $(\tilde{\mathcal{C}},\tilde{R})$ and the marked variables. For a more concise statement, we define an admissible set of initial variables, which is a set of variables that satisfy the conditions of the Starfish lemma.
Definition 4.8.
  1. 1)
    All variables in $A$ are regular;
  2. 2)
    All cluster variables in $A$ are pairwise coprime;
  3. 3)
Likewise, one defines an admissible set of initial variables $A\subseteq\tilde{\mathbf{x}}_{0}$ in $(\tilde{\mathcal{C}},\tilde{R})$. For the main result, we repeat the setup of the section.
Proposition 4.9.
@@ 4.3 Starfish lemma
    Each nonmarked cluster variable in $\mathbf{x}_{0}$ is coprime with each marked variable.
Then $\mathbf{x}_{0}$ is admissible, $\bar{\mathcal{A}}(\tilde{\mathcal{C}})\subseteq\tilde{R}$ and $\bar{\mathcal{A}}(\mathcal{C})\subseteq R$. If in addition
  1. 5)
    The exchange matrices $B_{0}$ and $\tilde{B}_{0}$ are of full rank;
  2. 6)
    There are inclusions $\mathcal{L}(\tilde{\mathbf{x}}_{0})\supseteq\tilde{R}$ and $\mathcal{L}(\tilde{\mathbf{x}}_{i}^{\prime})\supseteq\tilde{R}$ for each $i\in[1,\tilde{N}]$;
  3. 7)
    For each marked index $i\in\mathcal{I}(\kappa)$, $\mathcal{L}(\mathbf{x}_{i}^{\prime})\supseteq R$.
Then $\bar{\mathcal{A}}(\tilde{\mathcal{C}})=\tilde{R}$ and $\bar{\mathcal{A}}(\mathcal{C})=R$.
@@
Condition 2 constitutes conditions of the Starfish lemma 2.9, and hence $\bar{\mathcal{A}}(\tilde{\mathcal{C}})\subseteq\tilde{R}$. Conditions 5-6 allow us to apply Proposition 2.7 and conclude that $\bar{\mathcal{A}}(\tilde{\mathcal{C}})=\tilde{R}$.
To deduce $\bar{\mathcal{A}}(\mathcal{C})\subseteq R$, by the Starfish lemma, we need to show that $\mathbf{x}_{0}$ is admissible. By Proposition 4.4 and Condition 3, we see that all the initial variables in $\mathbf{x}_{0}$ are coprime; by Proposition 4.5 and Condition 4, we see that for each initial cluster variable $x\in\mathbf{x}_{0}$, $x^{\prime}$ is regular and coprime with $x$. We conclude that $\bar{\mathcal{A}}(\mathcal{C})\subseteq R$.
Lastly, by Proposition 5.3 and Condition 7, we conclude that $\bar{\mathcal{A}}(\mathcal{C})=R$. ∎
Remark 4.10.
In the proofs of Propositions 4.44.5 and 4.6, we didn’t use the specific form of the mutation relation, but only the following facts: 1) Starfish lemma; 2) the implication that if $\mathbf{x}_{0}$ is regular, then $x^{\prime}\in R[x^{\pm 1}]$ for every cluster variable $x\in\mathbf{x}_{0}$; 3) the equality $\bar{\mathcal{A}}(\mathcal{C})=\bar{\mathcal{A}}(\mathcal{N})$ for a star-shaped nerve $\mathcal{N}$ in $\mathcal{C}$; thus, Proposition 4.9 holds for more general cluster structures in which all the three conditions are fulfilled. In particular, if Remark 3.10 is taken into account, Proposition 4.9 holds for generalized cluster structures defined in [3, 15]. The same applies for Proposition 5.4 below.
4.4 Prime elements
Lemma 4.11.
Let $x\in R$ be a nonzero element that has only one minimal prime ideal $\mathfrak{p}$, and let $A$ be a multiplicatively closed subset of $R\setminus\{0\}$ such that $A\cap\mathfrak{p}=\emptyset$. If $x\cdot A^{-1}R=\mathfrak{p}\cdot A^{-1}R$, then $(x)=\mathfrak{p}$.
Proof.
Proposition 4.12.
Proof.
By Lemma 4.3, there is only one prime ideal $\mathfrak{p}$ of height $1$ such that $(x)\subseteq\mathfrak{p}$. Since $\mathcal{Q}$ is an isomorphism and $\tilde{x}$ is prime in $\tilde{R}$, $x$ is a prime element in $R[\mathbf{x}_{0}(\kappa)^{\pm 1}]$, and since $x$ is coprime with the marked variables, by Lemma 4.11, $(x)=\mathfrak{p}$. Thus $x$ is a prime element in $R$. ∎
Remark 4.13.
When a cluster structure is compatible with a Poisson bracket (in the sense of [12]), frozen variables are observed to have the following property: their zero loci foliate into unions of symplectic leaves of the ambient Poisson variety. To prove the latter, one may use an algebraic criterion from [18, Proposition 2.3] or [19, Remark 2.4], which in particular requires showing that a given frozen variable is a prime element in the coordinate ring of the given variety. Proposition 4.12 provides a way of transferring this property via a quasi-isomorphism. For an example of the use of Proposition 4.12 in this context, one may refer to our latest work [17].
4.5 Starfish lemma in factorial domains
In this subsection, we reformulate Proposition 4.9 in the case of factorial [^7] By a factorial domain we mean a Unique Factorization Domain. $\mathbb{C}$-algebras. Instead of assuming coprimality between the initial variables in $(\tilde{\mathcal{C}},\tilde{R})$, we assume that the initial variables are irreducible, and then apply Proposition 4.12 to derive irreducibility of the initial variables in $(\mathcal{C},R)$.
Definition 4.14.
  1. 1)
    All variables in $A$ are regular and irreducible;
  2. 2)
Proposition 4.15.
Let $(\mathcal{C},R)$ and $(\tilde{\mathcal{C}},\tilde{R})$ be cluster structures of geometric type in factorial Noetherian $\mathbb{C}$-algebras with no zero divisors and such that $R^{\times}=\tilde{R}^{\times}=\mathbb{C}^{\times}$. Let $(\mathbf{x}_{0},B_{0})$ and $(\tilde{\mathbf{x}}_{0},\tilde{B}_{0})$ be fixed initial seeds. Assume the following:
@@ 4.5 Starfish lemma in factorial domains
    The set of marked variables $\mathbf{x}_{0}(\kappa)$ is i-admissible.
Then $\mathbf{x}_{0}$ is i-admissible, as well as $\bar{\mathcal{A}}_{\mathbb{C}}(\tilde{\mathcal{C}})\subseteq\tilde{R}$ and $\bar{\mathcal{A}}_{\mathbb{C}}(\mathcal{C})\subseteq R$. If in addition
@@
    The exchange matrices $B_{0}$ and $\tilde{B}_{0}$ are of full rank;
  2. 5)
    There are inclusions $\mathcal{L}_{\mathbb{C}}(\tilde{\mathbf{x}}_{0})\supseteq\tilde{R}$ and $\mathcal{L}_{\mathbb{C}}(\tilde{\mathbf{x}}_{i}^{\prime})\supseteq\tilde{R}$ for each $i\in[1,\tilde{N}]$;
  3. 6)
    For each marked index $i\in\mathcal{I}(\kappa)$, $\mathcal{L}_{\mathbb{C}}(\mathbf{x}_{i}^{\prime})\supseteq R$.
Then $\bar{\mathcal{A}}_{\mathbb{C}}(\tilde{\mathcal{C}})=\tilde{R}$ and $\bar{\mathcal{A}}_{\mathbb{C}}(\mathcal{C})=R$.
@@
Since the initial variables $\tilde{\mathbf{x}}_{0}$ are algebraically independent over $\mathbb{Q}$, they are algebraically independent over $\mathbb{C}$, and since $\tilde{R}^{\times}=\mathbb{C}^{\times}$, the irreducibility of the initial variables implies their coprimality; in other words, $\tilde{\mathbf{x}}_{0}$ is admissible. By Proposition 4.12 and Condition 3, we see that $\mathbf{x}_{0}$ is i-admissible, and since the variables in $\mathbf{x}_{0}$ are algebraically independent and $R^{\times}=\mathbb{C}^{\times}$, $\mathbf{x}_{0}$ is admissible. Now the statement follows from Proposition 4.9. ∎
Remark 4.16.
The coordinate ring $\mathbb{C}[G]$ of a connected simply connected simple complex algebraic group $G$ is a factorial domain such that $(\mathbb{C}[G])^{\times}=\mathbb{C}^{\times}$ (see, for instance, [20]). Thus Proposition 4.15 covers the needs of the Gekhtman-Shapiro-Vainshtein program [11, 14] on the construction of multiple generalized cluster structures in $\mathbb{C}[G]$ for a given group $G$. At the cost of conciseness, Proposition 4.15 can be extended in various ways. For instance, one can require that $R^{\times}$ and $\tilde{R}^{\times}$ are groups of Laurent monomials in the invertible nonmarked frozen variables with coefficients in $\mathbb{C}^{\times}$ (thus covering the case of $\mathbb{C}[\operatorname{GL}_{n}]$). The proposition also extends to an arbitrary field extension of $\mathbb{Q}$; for fields of positive characteristic, one needs to additionally require that the initial variables remain algebraically independent over the new field.
@@ 5 Complementary birational quasi-isomorphisms
The objective of this section is to derive a result similar to Proposition 4.9 in the case when a cluster structure $(\mathcal{C},R)$ is related and birationally quasi-isomorphic to a pair of cluster structures. The existence of a pair of complementary birational quasi-isomorphisms allows us to impose fewer assumptions on the marked variables in $(\mathcal{C},R)$. The main result of this section is Proposition 5.4. The result is further refined in the case of factorial Noetherian domains in Proposition 5.8.
@@
5.1 Starfish lemma
In this subsection, we derive our second main result, which is Proposition 5.4.
@@ 5.1 Starfish lemma
    Each $x_{i}\in\mathbf{x}_{0}(\kappa)$ is coprime with each $x_{j}\in\mathbf{x}_{0}(\kappa^{\prime})$.
Then $x$ is regular and coprime with all the marked variables. If in addition $x$ is a cluster variable, $\tilde{x}^{\prime}$ is regular and coprime with $\tilde{x}$, and $\hat{x}^{\prime}$ is regular and coprime with $\hat{x}$, then $x^{\prime}$ is regular and coprime with $x$.
@@
Since $\mathcal{Q}$ induces an isomorphism of localizations $\mathcal{Q}:R[\mathbf{x}_{0}(\kappa)^{\pm 1}]\xrightarrow{\sim}\tilde{R}[\tilde{\mathbf{x}}_{0}(\kappa)^{\pm 1}]$ and since $\tilde{x}\in\tilde{R}$, we see that $x\in R[\mathbf{x}_{0}(\kappa)^{\pm 1}]$; likewise, since $\hat{x}$ is regular, we see that $x\in R[\mathbf{x}_{0}(\kappa^{\prime})^{\pm 1}]$. From Lemma 4.2 and Condition 3, we conclude that $R=R[\mathbf{x}_{0}(\kappa)^{\pm 1}]\cap R[\mathbf{x}_{0}(\kappa^{\prime})^{\pm 1}]$, and therefore, $x\in R$. To deduce coprimality between $x$ and a marked variable $x_{j}\in\mathbf{x}_{0}(\kappa^{\prime})$, we apply Proposition 4.4 relative $(\mathcal{C},\tilde{\mathcal{C}},\kappa)$; specifically, the variable $z$ from Proposition 4.4 is the variable $x$, and the variable $x$ from Proposition 4.4 is the variable $x_{j}$ (for which the coprimality with $\mathbf{x}_{0}(\kappa)$ is assumed); hence, $x$ is coprime with all the variable in $\mathbf{x}_{0}(\kappa^{\prime})$. Similarly, the coprimality between $x$ and the marked variables in $\mathbf{x}_{0}(\kappa)$ follows from Proposition 4.4 applied to $(\mathcal{C},\hat{\mathcal{C}},\kappa^{\prime})$. The last part follows from Proposition 4.5.∎
Proposition 5.3.
Assume that the initial exchange matrices in $\mathcal{C}$, $\tilde{\mathcal{C}}$ and $\hat{\mathcal{C}}$ are of full rank, as well as $\bar{\mathcal{A}}(\tilde{\mathcal{C}})\supseteq\tilde{R}$ and $\bar{\mathcal{A}}(\hat{\mathcal{C}})\supseteq\hat{R}$. Then $\bar{\mathcal{A}}(\mathcal{C})\supseteq R$.
Proof.
Now we formulate our second main result. Recall that a subset of initial variables is admissible if it satisfies the conditions of the Starfish lemma (see Definition 4.8).
Proposition 5.4.
Let $R$, $\tilde{R}$ and $\hat{R}$ be normal Noetherian domains, $(\mathcal{C},R)$, $(\tilde{\mathcal{C}},\tilde{R})$ and $(\hat{\mathcal{C}},\hat{R})$ be cluster structures of geometric type in $R$, $\tilde{R}$ and $\hat{R}$, with fixed initial extended seeds $\Sigma_{0}:=(\mathbf{x}_{0},B_{0})$, $\tilde{\Sigma}_{0}:=(\tilde{\mathbf{x}}_{0},\tilde{B}_{0})$ and $\hat{\Sigma}_{0}:=(\hat{\mathbf{x}}_{0},\hat{B}_{0})$. Assume the following:
@@
    There are markers $\kappa$ and $\kappa^{\prime}$ that make $(\mathcal{C},\tilde{\mathcal{C}},\kappa)$ and $(\mathcal{C},\hat{\mathcal{C}},\kappa^{\prime})$ related triples, as well as a pair of birational quasi-isomorphisms $\mathcal{Q}:(\mathcal{C},R)\rightarrow(\tilde{\mathcal{C}},\tilde{R})$ and $\mathcal{Q}^{\prime}:(\mathcal{C},R)\rightarrow(\hat{\mathcal{C}},\hat{R})$;
@@
    Each $x_{i}\in\mathbf{x}_{0}(\kappa)$ is coprime with each $x_{j}\in\mathbf{x}_{0}(\kappa^{\prime})$.
Then $\mathbf{x}_{0}$ is admissible, as well as $\bar{\mathcal{A}}(\hat{\mathcal{C}})\subseteq\hat{R}$, $\bar{\mathcal{A}}(\tilde{\mathcal{C}})\subseteq\tilde{R}$ and $\bar{\mathcal{A}}(\mathcal{C})\subseteq R$. If in addition,
@@
    The initial exchange matrices $B_{0}$, $\tilde{B}_{0}$ and $\hat{B}_{0}$ are of full rank;
  2. 5)
    For the star-shaped nerves $\tilde{\mathcal{N}}$ and $\hat{\mathcal{N}}$ centered at $\tilde{\Sigma}_{0}$ and $\hat{\Sigma}_{0}$, $\bar{\mathcal{A}}(\tilde{\mathcal{N}})\supseteq\tilde{R}$ and $\bar{\mathcal{A}}(\hat{\mathcal{N}})\supseteq\hat{R}$.
Then $\bar{\mathcal{A}}(\hat{\mathcal{C}})=\hat{R}$, $\bar{\mathcal{A}}(\tilde{\mathcal{C}})=\tilde{R}$ and $\bar{\mathcal{A}}(\mathcal{C})=R$.
@@
As in the proof of Proposition 4.9, Condition 2 is enough to conclude $\bar{\mathcal{A}}(\hat{\mathcal{C}})\subseteq\hat{R}$ and $\bar{\mathcal{A}}(\tilde{\mathcal{C}})\subseteq\tilde{R}$. To claim $\bar{\mathcal{A}}(\mathcal{C})\subseteq R$, we need to show that $\mathbf{x}_{0}$ is admissible. From Conditions 2, 3 and Proposition 5.2, we see that each initial variable $x\in\mathbf{x}_{0}$ is coprime with all the marked variables, and if $x$ is a cluster variable, the last part of Proposition 5.2 shows that $x^{\prime}$ is regular and coprime with $x$. By Proposition 4.4 applied to $x$ and any other initial variable $z$, we conclude that $x$ and $z$ are coprime. Therefore, $\mathbf{x}_{0}$ is admissible, and thus $\bar{\mathcal{A}}(\mathcal{C})\subseteq R$.
For the second part, Conditions 4 and 5 together with Proposition 2.7 imply $\bar{\mathcal{A}}(\hat{\mathcal{C}})=\hat{R}$ and $\bar{\mathcal{A}}(\tilde{\mathcal{C}})=\tilde{R}$; by Proposition 5.3, we conclude that $\bar{\mathcal{A}}(\mathcal{C})=R$. ∎
5.2 Prime elements
We continue in the setup of the section. Here, we show that if $\tilde{x}$ and $\hat{x}$ are variables in $\tilde{\mathcal{C}}$ and $\hat{\mathcal{C}}$ that are regular and prime, then, under certain conditions, the related variable $x$ is a prime element in $R$. Some applications of this type of results are outlined in Remark 4.13.
Proposition 5.5.
Let $x$ be a cluster or frozen variable (marked or nonmarked) from an extended marked cluster in $\mathcal{C}$. Assume the following:
  1. 1)
  2. 2)
  3. 3)
    Each $x_{i}\in\mathbf{x}_{0}(\kappa)$ is coprime with each $x_{j}\in\mathbf{x}_{0}(\kappa^{\prime})$;
  4. 4)
    The variable $\tilde{x}$ is prime in $\tilde{R}$ or $\hat{x}$ is prime in $\hat{R}$.
  5. 5)
    The variable $x$ is not invertible in $R$.
Then $x$ is a prime element in $R$.
Proof.
The first three conditions are exactly the conditions in Proposition 5.2, and therefore, $x$ is regular and coprime with all the marked variables. By Condition 4, without loss of generality, assume that $\tilde{x}$ is prime in $\tilde{R}$. Since $\mathcal{Q}$ induces a bijection $\mathcal{P}_{\kappa}(\tilde{x})\cong\mathcal{P}_{\kappa}(x)=\mathcal{P}(x)$ (see Lemma 4.3) and $x$ is not invertible by Condition 5, we see that $|\mathcal{P}_{\kappa}(\tilde{x})|=1$. Since $\tilde{x}$ is itself prime, the latter implies $\tilde{x}$ is coprime with $\tilde{\mathbf{x}}_{0}(\kappa)$. It follows from Proposition 4.12 that $x$ is prime in $R$. ∎
5.3 Starfish lemma in factorial domains
We continue in the setup of the previous subsections. To remind the reader, our objective was to derive the conditions of the Starfish lemma for $(\mathcal{C},R)$ assuming that they hold for $(\tilde{\mathcal{C}},\tilde{R})$ and $(\hat{\mathcal{C}},\hat{R})$. For our applications, we would like to impose as few assumptions on the marked variables as possible. Comparing Proposition 5.4 with Proposition 4.9, we see that the existence of a pair of complementary birational quasi-isomorphisms reduces the number of assumptions for the marked variables (yet, it does not eliminate these assumptions completely). The only assumption on the marked variables in $(\mathcal{C},R)$ is:
$\text{Each}\ x_{i}\in\mathbf{x}_{0}(\kappa)\ \text{is coprime with each}\ x_{j}\in\mathbf{x}_{0}(\kappa^{\prime}).$ (5.3)
Similar to Proposition 4.15, in this subsection we impose extra conditions on the rings $R$, $\tilde{R}$ and $\hat{R}$, as well as require the initial variables in $\tilde{\mathcal{C}}$ and $\hat{\mathcal{C}}$ to be irreducible. As one shall see, this allows us to drop Condition 5.3. Let us introduce the following notation:

### v3

@@ 5.2 Transferring cluster and algebraic properties
In this subsection, we show how to derive coprimality and regularity of nonmarked variables in $\mathcal{C}$ given their regularity and coprimality in $\tilde{\mathcal{C}}$.
@@
    Both $\tilde{x}$ and its mutation $\tilde{x}^{\prime}$ are regular and coprime;
  2. 2)
@@
  3. 3)
    The exchange polynomial $xx^{\prime}$ is regular.
@@
5.3 Prime elements
Lemma 5.5.
Let $x\in R$ be an element that has only one minimal prime ideal $\mathfrak{p}$, and let $A$ be a multiplicatively closed subset of $R\setminus\{0\}$ such that $A\cap\mathfrak{p}=\emptyset$. If $x\cdot A^{-1}R=\mathfrak{p}\cdot A^{-1}R$, then $(x)=\mathfrak{p}$.
Proof.
@@ 5.3 Prime elements
By Lemma 5.2, there is only one prime ideal $\mathfrak{p}$ of height $1$ such that $(x)\subseteq\mathfrak{p}$. Since $\mathcal{Q}$ is an isomorphism and $\tilde{x}$ is prime in $\tilde{R}$, $x$ is a prime element in $R[\mathbf{x}_{0}(\kappa)^{\pm 1}]$; hence, by Lemma 5.5, $(x)=\mathfrak{p}$. Thus $x$ is a prime element in $R$. ∎
@@
When a cluster structure is compatible with a Poisson bracket (in the sense of [12]), frozen variables are observed to have the following property: their zero loci foliate into unions of symplectic leaves of the ambient Poisson variety. To prove the latter, one may use an algebraic criterion from [18, Proposition 2.3] or [20, Remark 2.4], which in particular requires showing that a given frozen variable is a prime element in the coordinate ring of the given variety. Proposition 5.6 provides a way of transferring this property via a birational quasi-isomorphism. For an example of the use of Proposition 5.6 in this context, one may refer to our latest work [17].
5.4 Proof of Proposition 4.2
Recall that a set of initial variables is admissible if it satisfies the conditions of the Starfish lemma (see Definition 4.1). In this subsection, we prove Proposition 4.2, which we restate for convenience.
Proposition.
Let $(\mathcal{C},R)$ and $(\tilde{\mathcal{C}},\tilde{R})$ be cluster structures of geometric type in normal Noetherian domains $R$ and $\tilde{R}$ with fixed initial extended clusters $\mathbf{x}_{0}$ and $\tilde{\mathbf{x}}_{0}$. Assume the following:
@@
    Each nonmarked cluster variable in $\mathbf{x}_{0}$ is regular and is coprime with each marked variable.
Then $\mathbf{x}_{0}$ is admissible, $\bar{\mathcal{A}}(\tilde{\mathcal{C}})\subseteq\tilde{R}$ and $\bar{\mathcal{A}}(\mathcal{C})\subseteq R$.
@@
It follows from the Starfish lemma and Condition 2 that $\bar{\mathcal{A}}(\tilde{\mathcal{C}})\subseteq\tilde{R}$. To deduce $\bar{\mathcal{A}}(\mathcal{C})\subseteq R$, we need to show that $\mathbf{x}_{0}$ is admissible. By Proposition 5.3 and Condition 3, we see that the initial cluster variables in $\mathbf{x}_{0}$ are coprime; by Proposition 5.4 and Condition 4, we see that for each initial cluster variable $x\in\mathbf{x}_{0}$, $x^{\prime}$ is regular and coprime with $x$. Thus $\bar{\mathcal{A}}(\mathcal{C})\subseteq R$. ∎
5.5 Proof of Proposition 4.5
Recall that an initial set of variables is i-admissible if the variables are regular, irreducible, and (for cluster variables) coprime with their mutations (see Definition 4.4). In this subsection, we prove Proposition 4.5, which we restate for convenience.
Proposition.
Let $(\mathcal{C},R)$ and $(\tilde{\mathcal{C}},\tilde{R})$ be cluster structures of geometric type in factorial Noetherian $\mathbb{C}$-algebras with no zero divisors and such that $R^{\times}=\tilde{R}^{\times}=\mathbb{C}^{\times}$. Let $\mathbf{x}_{0}$ and $\tilde{\mathbf{x}}_{0}$ be fixed initial extended clusters. Assume the following:
@@
    The set of marked variables $\mathbf{x}_{0}(\kappa)$ is i-admissible;
@@
    Each nonmarked cluster variable in $\mathbf{x}_{0}$ is regular and is coprime with each marked variable.
Then $\mathbf{x}_{0}$ is i-admissible, as well as $\bar{\mathcal{A}}_{\mathbb{C}}(\tilde{\mathcal{C}})\subseteq\tilde{R}$ and $\bar{\mathcal{A}}_{\mathbb{C}}(\mathcal{C})\subseteq R$.
@@
Since the initial variables $\tilde{\mathbf{x}}_{0}$ are algebraically independent over $\mathbb{Q}$, they are algebraically independent over $\mathbb{C}$, and since $\tilde{R}^{\times}=\mathbb{C}^{\times}$, the irreducibility of the initial variables implies their coprimality; in other words, $\tilde{\mathbf{x}}_{0}$ is admissible. Likewise, $\mathbf{x}_{0}(\kappa)$ is admissible. By Proposition 5.6 and Conditions 3-4, we see that $\mathbf{x}_{0}$ is i-admissible, and since the variables in $\mathbf{x}_{0}$ are algebraically independent and $R^{\times}=\mathbb{C}^{\times}$, $\mathbf{x}_{0}$ is admissible. Now the statement follows from Proposition 4.2. ∎
5.6 Proof of Proposition 4.9
We restate Proposition 4.9 for convenience.
Proposition.
  1. 1)
  2. 2)
    The matrices $B_{0}$ and $\tilde{B}_{0}$ are of full rank;
  3. 3)
    The reverse inclusion $\bar{\mathcal{A}}(\tilde{\mathcal{C}})\supseteq\tilde{R}$ holds;
  4. 4)
    For each $i\in\mathcal{I}(\kappa)$, $\mathcal{L}(\mathbf{x}_{i}^{\prime})\supseteq R$.
Then $\bar{\mathcal{A}}(\mathcal{C})\supseteq R$.
Proof.
@@ 6 Complementary birational quasi-isomorphisms
In this section, we prove the results that involve the use of a pair of complementary birational quasi-isomorphisms; that is, Proposition 4.3, Proposition 4.5 and Proposition 4.11.
@@
6.1 Transferring cluster and algebraic properties
We begin with an auxiliary result.
@@ 6.1 Transferring cluster and algebraic properties
    Every $x_{i}\in\mathbf{x}_{0}(\kappa)$ is coprime with every $x_{j}\in\mathbf{x}_{0}(\kappa^{\prime})$.
Then $x$ is regular and coprime with all the marked variables.
@@
Since $\mathcal{Q}$ induces an isomorphism of localizations $\mathcal{Q}:R[\mathbf{x}_{0}(\kappa)^{\pm 1}]\xrightarrow{\sim}\tilde{R}[\tilde{\mathbf{x}}_{0}(\kappa)^{\pm 1}]$ and since $\tilde{x}\in\tilde{R}$, we see that $x\in R[\mathbf{x}_{0}(\kappa)^{\pm 1}]$; likewise, since $\hat{x}$ is regular, we see that $x\in R[\mathbf{x}_{0}(\kappa^{\prime})^{\pm 1}]$. From Lemma 5.1 and Condition 3, we conclude that $R=R[\mathbf{x}_{0}(\kappa)^{\pm 1}]\cap R[\mathbf{x}_{0}(\kappa^{\prime})^{\pm 1}]$, and therefore, $x\in R$. To deduce coprimality between $x$ and a marked variable $x_{j}\in\mathbf{x}_{0}(\kappa^{\prime})$, we apply Proposition 5.3 relative $(\mathcal{C},\tilde{\mathcal{C}},\kappa)$; specifically, the variable $z$ from Proposition 5.3 is the variable $x$, and the variable $x$ from Proposition 5.3 is the variable $x_{j}$ (for which the coprimality with $\mathbf{x}_{0}(\kappa)$ is assumed); hence, $x$ is coprime with all the variables in $\mathbf{x}_{0}(\kappa^{\prime})$. Similarly, the coprimality between $x$ and the marked variables in $\mathbf{x}_{0}(\kappa)$ follows from Proposition 5.3 applied to $(\mathcal{C},\hat{\mathcal{C}},\kappa^{\prime})$. ∎
6.2 Proof of Proposition 4.3
We restate Proposition 4.3 for convenience.
Proposition.
Let $R$, $\tilde{R}$ and $\hat{R}$ be normal Noetherian domains, $(\mathcal{C},R)$, $(\tilde{\mathcal{C}},\tilde{R})$ and $(\hat{\mathcal{C}},\hat{R})$ be cluster structures of geometric type in $R$, $\tilde{R}$ and $\hat{R}$, with fixed initial extended clusters $\mathbf{x}_{0}$, $\tilde{\mathbf{x}}_{0}$ and $\hat{\mathbf{x}}_{0}$. Assume the following:
@@
    There are markers $\kappa$ and $\kappa^{\prime}$ that make $(\mathcal{C},\tilde{\mathcal{C}},\kappa)$ and $(\mathcal{C},\hat{\mathcal{C}},\kappa^{\prime})$ related triples, as well as a pair of complementary birational quasi-isomorphisms $\mathcal{Q}:(\mathcal{C},R)\rightarrow(\tilde{\mathcal{C}},\tilde{R})$ and $\mathcal{Q}^{\prime}:(\mathcal{C},R)\rightarrow(\hat{\mathcal{C}},\hat{R})$;
@@
    Every $x_{i}\in\mathbf{x}_{0}(\kappa)$ is coprime with every $x_{j}\in\mathbf{x}_{0}(\kappa^{\prime})$;
@@
    The initial extended cluster $\mathbf{x}_{0}$ is regular.
Then $\mathbf{x}_{0}$ is admissible, as well as $\bar{\mathcal{A}}(\hat{\mathcal{C}})\subseteq\hat{R}$, $\bar{\mathcal{A}}(\tilde{\mathcal{C}})\subseteq\tilde{R}$ and $\bar{\mathcal{A}}(\mathcal{C})\subseteq R$.
@@
As in the proof of Proposition 4.2, Condition 2 is enough to conclude $\bar{\mathcal{A}}(\hat{\mathcal{C}})\subseteq\hat{R}$ and $\bar{\mathcal{A}}(\tilde{\mathcal{C}})\subseteq\tilde{R}$. To claim $\bar{\mathcal{A}}(\mathcal{C})\subseteq R$, we need to show that $\mathbf{x}_{0}$ is admissible. From Conditions 2, 3 and Proposition 6.2, we see that each nonmarked initial variable $x\in\mathbf{x}_{0}$ is coprime with all the marked variables, and if $x$ is a cluster variable, by Proposition 5.4, $x^{\prime}$ is regular and coprime with $x$. By Proposition 5.3 applied to $x,z\in\mathbf{x}_{0}\setminus\mathbf{x}_{0}(\kappa)$ relative $(\mathcal{C},\tilde{\mathcal{C}},\kappa)$, we see that $x$ and $z$ are coprime (and likewise, if $x,z\in\mathbf{x}\setminus\mathbf{x}_{0}(\kappa^{\prime})$, apply the proposition relative $(\mathcal{C},\hat{\mathcal{C}},\kappa^{\prime})$). By Proposition 5.4, the mutation $x^{\prime}$ of a marked variable $x$ is regular and coprime with $x$. Therefore, $\mathbf{x}_{0}$ is admissible, and thus $\bar{\mathcal{A}}(\mathcal{C})\subseteq R$. ∎
6.3 Proof of Proposition 4.6
For this subsection, let us introduce additional notation: